OpenStudy (anonymous):

f(x) = -2x-1 and h(x) = 3/x+1 what does (h*f)(2) = ??

hero (hero):

f(x) = -2x-1 h(x) = 3/(x+1) f(2) = -2(2) - 1 = -4 - 1 = -5 h(2) = 3/(2+1) = 3/3 = 1 (h*f)(2) = h(2)*f(2) = 1*-5 = -5

OpenStudy (mertsj):

f(2)=-5 h(2)=1 f(2)*h(2)=-5(1)=-5

OpenStudy (anonymous):

Is that a valid way of writing f(2)*h(2) ?

hero (hero):

* means times. It's valid for when you're using a computer or a calculator

OpenStudy (anonymous):

I meant (h*f) seems like multiplying "functions" rather than their values. Do you understand what I mean?

hero (hero):

Yes, you can also multiply the functions, but that would take longer. It's easier to evaluate f(2) and h(2) then multiply that together. Either way, you would get the same answer.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

hmmmm......i got the answer -3/4 ?? that is the answer in the book i am using.

OpenStudy (anonymous):

depends on whether you want \[f\times g(x)\] or \[f\circ g(x)\] they are different

OpenStudy (anonymous):

ok that makes sense b/c the question was (h o f)(2)

OpenStudy (anonymous):

then \[h\circ f(2)=h(f(2))=h(-5)=\frac{3}{-5+1}=-\frac{3}{4}\]